Having a hard time figuring out this equation in a simplified fraction form.
0.19x+0.2(x-4)=0.01(3x-5)

Answer by fractalier(6550)

The easiest way out of these is to multiply everything by the lowest common denominator...here that is 100...so from
0.19x+0.2(x-4)=0.01(3x-5)
we multiply by 100 and get
19x + 20(x-4) = 3x - 5
19x + 20x - 80 = 3x - 5
39x - 80 = 3x -5
36x = 75
x = 75/36 = 25/12